Show that the adjusted sample variance is a consistent estimator of the true variance


1
Expert's answer
2021-08-10T11:12:52-0400

Solution:

Let "X_{1}, X_{2} \ldots \ldots \ldots X_{n}" be the independent observations from a population with mean "\mu" and variance "\sigma^{2}"

"E\left(X_{i}\right)=\mu, \operatorname{Var}\left(X_{i}\right)=\sigma^{2}

\\E\left(X^{2}\right)=\sigma^{2}+\mu^{2}

\\\operatorname{Var}(X)=E\left(X^{2}\right)-[E(X)]^{2}

\\E\left(\bar{X}^{2}\right)=\frac{\sigma^{2}}{n}+\mu^{2}"

Now, we show that

"{E}\left(s^{2}\right)={E}\left(\frac{\sum_{i=1}^{n}\left(X_{i}-\bar{X}\right)^{2}}{n-1}\right)=\sigma^{2}"

"\sum" is going from 1 to n

"{E}\left(\sum\left(X_{i}-\bar{X}\right)^{2}\right)={E}\left(\sum X_{i}^{2}-2 \bar{X} \sum X_{i}+n \bar{X}^{2}\right)=\sum {E}\left(X_{i}^{2}\right)-{E}\left(n \bar{X}^{2}\right)

\\\sum {E}\left(X_{i}^{2}\right)-{E}\left(n \bar{X}^{2}\right)=\sum {E}\left(X_{i}^{2}\right)-n {E}\left(\bar{X}^{2}\right)=n \sigma^{2}+n \mu^{2}-\sigma^{2}-n \mu^{2}"

It will be simplified to "(\mathrm{n}-1) \sigma^{2}"

Therefore,

"{E}\left(\sum\left(X_{i}-\bar{X}\right)^{2}\right)=(n-1) \sigma^{2}
\\{E}\left(s^{2}\right)={E}\left(\frac{\Sigma\left(X_{i}-\bar{X}\right)^{2}}{n-1}\right)=\frac{1}{n-1} {E}\left(\sum\left(X_{i}-\bar{X}\right)^{2}\right)"

"{E}(s^2)=\dfrac{(n-1) \sigma^{2}}{n-1}=\sigma^{2}"

Hence, it is proved that that the sample variance is an unbiased estimator of the population variance. Replacing X with the Y

"E[\dfrac1{n-1}\sum_{i=1}^{n}\left(Y_{i}-\bar{Y}\right)]={\sigma}^2"

Hence, proved.